# VideoQA-SC: Adaptive Semantic Communication for Video Question Answering

### v1

@@ 2.1 Spatiotemporal Semantic Encoder
To reduce the use of computational and transmission resources, we initially apply uniform interval sparse sampling across each video, choosing $l_{v}$ frames to form the keyframes that make up $\mathbf{X}_{v}$. The $l_{v}$ keyframes are divided into $l_{c}$ clips, each with a length of $l_{f}$ frames. For simplicity, we assume that each clip operates independently, and $g(\cdot;\boldsymbol{\zeta})$ is designed to only capture the correlations within the frames of a single clip and produce semantics for the clip.
The process of temporal and spatial modeling is illustrated in Fig. 3. For the $i$-th clip $\mathcal{X}_{i}=\left\{\mathbf{X}_{i,1},\mathbf{X}_{i,2},\cdots,\mathbf{X}_{i,l_{f}}\right\}$, $\mathbf{X}_{i,j}\in\mathbb{R}^{3\times{x}\times{y}}$ represents the $j$-th frame in the clip. We first use a pre-trained object detector and ResNet to process all $\mathbf{X}_{i,j}$ $(j\in{\left[1,l_{f}\right]})$, obtaining the object-level feature $\mathbf{O}_{i,j}\in{\mathbb{R}^{r\times{m}}}$ and frame-level feature $\mathbf{F}_{i,j}\in{\mathbb{R}^{m}}$, respectively. Let $r$ be the number of detected objects in $\mathbf{X}_{i,j}$ and $m$ be the channel dimension of both the object-level features and the frame-level features. We concatenate all $\mathbf{O}_{i,j}$ and $\mathbf{F}_{i,j}$ to get the $i$-th clip object-level feature $\mathbf{O}_{i}\in{\mathbb{R}^{l_{f}\times{r}\times{m}}}$ and the $i$-th clip frame-level feature $\mathbf{F}_{i}\in{\mathbb{R}^{l_{f}\times{m}}}$, respectively.
Then, each $\mathbf{O}_{i}$ is fed into stacked Transformer blocks to facilitate the interaction of the same object features in different frames in the $i$-th clip. Here, the number of frames in the clip corresponds to the sequence length in the original Transformer. By using self-attention mechanism, we can obtain an aggregated representation of each object $\mathbf{O}_{i,j,k}\in\mathbb{R}^{m}$ $(k\in\left[1,r\right])$ within one clip to learn the object-level temporal correlations. The aggregation can be expressed as:
@@
where $\alpha_{i,k,j,p}$ is the attention score of aggregated $\mathbf{O}_{i,j,k}$ to $\mathbf{O}_{i,p,k}$.
Subsequently, following the work [27], we construct a graph $\mathcal{G}_{i}$ for every $\mathbf{O}_{i,j}$, where each object in $\mathbf{O}_{i,j}$ is a node in $\mathcal{G}_{i}$. Based on the constructed graph, we perform graph convolution operations on every $\mathbf{O}_{i,j}$, exploiting the structural information between different objects in one frame to utilize spatial interaction. Finally, after average pooling along the object dimension, the processed object-level feature $\mathbf{O^{\prime}}_{i}\in\mathbb{R}^{l_{f}\times{m}}$ is concatenated with the frame-level feature $\mathbf{F}_{i}$. The concatenated feature is fed into a linear layer to map to the video semantics $\mathbf{Y}_{v,i}\in\mathbb{R}^{l_{f}\times{d}}$. All clip-level representations $\mathbf{Y}_{v,i}$ are concatenated to obtain the entire video semantics $\mathbf{Y}_{v}\in\mathbb{R}^{l_{v}\times{d}}$.
@@
2.2 Cross-Attention Based DJSCC Transmission
@@
We use Transformer to construct the backbone of both the JSC encoder and decoder. In the Transformer architecture, the self-attention mechanism utilizes the learnable matrix $\mathbf{W}^{QKV}$ to generate the query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$) representations from embeddings. This process enables weighted information aggregation, effectively capturing dependencies between different embeddings in the sequence. In DJSCC, code rate guidance allows the refinement of the latent representations, thereby generating channel input symbols adapted to bandwidth constraints, which motivates us to provide code rate guidance to the encoding and decoding processes of video semantics.
Different from directly using the human-defined code rate, we design the code rate as learnable parameters and interact with the video semantics in the form of cross attention. Specifically, we introduce a learnable parameter tensor, with the same shape as $\mathbf{Y}_{v}$, termed rate embedding $\mathbf{Y}_{\text{rate}}$, in the JSC encoder $f_{e}(\cdot;\boldsymbol{\theta},\boldsymbol{\epsilon})$ and the JSC decoder $f_{d}(\cdot;\boldsymbol{\phi})$. The encoding and decoding processes of $\mathbf{Y}_{v}$ are both guided by $\mathbf{Y}_{\text{rate}}$ shared between the transmitter and receiver. Furthermore, $\mathbf{Y}_{\text{rate}}$ can aid the JSC encoder in achieving variable-length coding of $\mathbf{Y}_{v}$, which will be explained in detail in Section 2.3.
@@
$\tilde{\mathbf{Y}}_{\text{rate}}^{l}&=\operatorname{MHCA}(\operatorname{LN}(\mathbf{Y}_{\text{rate}}),\operatorname{LN}(\mathbf{Y}_{v}))+\mathbf{Y}_{\text{rate}}^{l-1},\\ \mathbf{Y}_{\text{rate}}^{l}&=\operatorname{FFN}(\operatorname{LN}(\tilde{\mathbf{Y}}_{\text{rate}}^{l}))+\operatorname{LN}(\tilde{\mathbf{Y}}_{\text{rate}}^{l}),\end{aligned}$ (10)
@@
$\tilde{\mathbf{Y}}_{v}^{l}&=\operatorname{MHCA}(\operatorname{LN}(\mathbf{Y}_{v}),\operatorname{LN}(\mathbf{Y}_{\text{rate}}))+\mathbf{Y}_{v}^{l-1},\\ \mathbf{Y}_{v}^{l}&=\operatorname{FFN}(\operatorname{LN}(\tilde{\mathbf{Y}}_{v}^{l}))+\operatorname{LN}(\tilde{\mathbf{Y}}_{v}^{l}),\end{aligned}$ (11)
respectively. In Eq. (10) and (11), $\mathbf{Y}_{\text{rate}}^{l-1}$ and $\mathbf{Y}_{v}^{l-1}$ denote the inputs of the $l$-th Transformer block of the two branches, and $\mathbf{Y}_{\text{rate}}^{l}$ and $\mathbf{Y}_{v}^{l}$ denote the outputs of the $l$-th Transformer blocks of the two branches. $\operatorname{MHCA}(\cdot)$ represents $\operatorname{CA}(\cdot)$ function with multi-head. $\operatorname{LN}(\cdot)$ represents the layer normalization in the Transformer. $\operatorname{FFN}(\cdot)$ represents two linear layers with $\operatorname{GeLU}(\cdot)$ as the activation function.
@@
2.3 Learning-Based Adaptive Bandwidth Allocation
@@
Figure 5: The JSC encoder with adaptive bandwidth allocation. CA Transformer block denotes the proposed dual-branch cross-attention Transformer block in Section 2.2.
According to the cross-attention mechanism, $\mathbf{Y}_{\text{rate}}$ described in Section Section 2.2 can be seen as the score metric that dynamically scales elements of $\mathbf{Y}_{v}$, causing us to measure the importance of features based on $\mathbf{Y}_{\text{rate}}$. Our learning-based adaptive bandwidth allocation approach exploits $\mathbf{Y}_{\text{rate}}$ as guidance, which sparsifies the channels of each token output by the JSC encoder to generate channel input symbols of variable lengths. Specifically, given the output of the last dual-branch Transformer block $\mathbf{Y}_{\text{rate}}\in{\mathbb{R}^{l_{v}\times d}}$ and $\mathbf{Y}_{v}\in{\mathbb{R}^{l_{v}\times d}}$, we develop a series of rate predictors parameterized by $\boldsymbol{\epsilon}$ to predict the retained dimension for each token $\mathbf{Y}_{v,i}\in{\mathbb{R}^{d}}$ by $\mathbf{Y}_{\text{rate}}$. Then, a binary mask matrix $\mathbf{M}\in{\mathbb{R}^{l_{v}\times d}}$ is generated for channel masking, with the $i$-th row containing the first $k_{i}$ ones followed by $(d-k_{i})$ zeros ($k_{i}\in{\left[0,d\right]}$). $\mathbf{M}$ is used to retain the first $k_{i}$ channels of $\mathbf{Y}_{v,i}$ and mask the rest channels:
@@
where $\operatorname{AveragePool}(\cdot)$ denotes the average pooling operation. After that, the rate predictor combines the local and global information along the channel dimension:
@@
At the receiver, we first generate $\mathbf{M}$ based on received $\mathbf{b}$. Then, we unflatten and zero-pad noisy real channel output symbols $\hat{\mathbf{s}}_{v}$ based on $\mathbf{M}$ to get $\hat{\mathbf{S}}_{v}$. The learnable vector $\mathbf{c}\in\mathbb{R}^{d}$ is developed to compensate for the information lost due to the channel masking operation. For each token $\hat{\mathbf{S}}_{v,i}\in\mathbb{R}^{d}$, if its $j$-th channel is masked, $\mathbf{c}_{j}$ is selected as the initial value for this channel:
@@
2.4 Content-Adaptive and SNR-Adaptive
The bandwidth allocation method described in Section 2.3 can achieve bandwidth efficiency for VideoQA-SC. The decision $\mathbf{D}$ is only determined by a series versions of $\mathbf{Y}_{\text{rate}}$ ($\mathbf{Y}_{\text{rate}}^{1}$,$\mathbf{Y}_{\text{rate}}^{2}$,…,$\mathbf{Y}_{\text{rate}}^{L}$) and a series versions of $\mathbf{Y}_{v}$ ($\mathbf{Y}_{v}^{1}$,$\mathbf{Y}_{v}^{2}$,…,$\mathbf{Y}_{v}^{L-1}$), which indicates that the bandwidth allocation is adaptive to video semantics or video contents. Such rate predictors will make the same bandwidth allocation under different channel conditions, which is inconsistent with traditional channel coding ideas. Since content-adaptive bandwidth allocation is not robust to noise, it is difficult to support SC under diverse channel conditions.
@@ 2.5 Multimodal Fuser
Given $\hat{\mathbf{Y}}_{v}$ and $\mathbf{Y}_{q}$, the interaction of the two modal information can be achieved through attention-based weighted fusion. After linear projection, $\hat{\mathbf{Y}}_{v}$ is mapped to $\mathbf{E}_{v}\in\mathbb{R}^{l_{v}\times d}$ as query and $\mathbf{Y}_{q}$ is mapped to $\mathbf{E}_{q}\in\mathbb{R}^{b\times{s_{i}}\times d}$ as key. Then, we caculate the attention of $\hat{\mathbf{Y}}_{v}$ to each candidate QA pair feature $\mathbf{Y}_{q,i}$, and utilize attention-based fusion of $\hat{\mathbf{Y}}_{v}$ and $\mathbf{Y}_{q}$ to get the QA-aware video feature $\mathbf{Y}_{qv}\in\mathbb{R}^{l_{v}\times d}$, which can be formulated as:
@@
$a^{\star}=\mathop{\arg\max}\limits_{i}{a_{\text{pred},i}},$ (27)
where ${a_{\text{pred},i}}$ is the $i$-th element of $\mathbf{a}_{\text{pred}}$.
@@
2.6 Training Strategy

### v2

@@ III-A Spatiotemporal Semantic Encoder
To reduce the use of computational and transmission resources, we initially apply uniform interval sparse sampling across each video, choosing $l_{v}$ frames as keyframes to make up $\mathbf{X}_{v}$. The $l_{v}$ keyframes are divided into $l_{c}$ clips, each with a length of $l_{f}$ frames. For simplicity, we assume that each clip operates independently, and $g(\cdot;\boldsymbol{\zeta})$ is designed to only capture the correlations within the frames of a single clip and produce semantics for the clip.
Figure 3: Temporal and spatial modeling of the $i$-th video clip. Shapes with the same color represent the same objects in different frames. Striped shapes indicate fused features after corresponding modeling. An object-level feature extractor and a frame-level feature extractor are used to preprocess all frames in the $i$-th clip for further processing.
The process of temporal and spatial modeling is illustrated in Fig. 3. For the $i$-th clip $\mathcal{X}_{i}=\left\{\mathbf{X}_{i,1},\mathbf{X}_{i,2},\cdots,\mathbf{X}_{i,l_{f}}\right\}$, $\mathbf{X}_{i,j}\in\mathbb{R}^{3\times{x}\times{y}}$ represents the $j$-th frame in the clip. We first use a pretrained object detector from [41] and pretrained standard ResNet-101 from [42] without classification heads as the preprocessing networks to process all $\mathbf{X}_{i,j}$ $(j\in{\left[1,l_{f}\right]})$, obtaining the object-level feature $\mathbf{O}_{i,j}\in{\mathbb{R}^{r\times{m}}}$ and frame-level feature $\mathbf{F}_{i,j}\in{\mathbb{R}^{m}}$, respectively. Actually, any network which is able to extract object-level or frame-level features can serve as the preprocessing network. Let $r$ be the number of detected objects in $\mathbf{X}_{i,j}$ and $m$ be the channel dimension of both the object-level features and the frame-level features. We concatenate all $\mathbf{O}_{i,j}$ and $\mathbf{F}_{i,j}$ to get the $i$-th clip object-level feature $\mathbf{O}_{i}\in{\mathbb{R}^{l_{f}\times{r}\times{m}}}$ and the $i$-th clip frame-level feature $\mathbf{F}_{i}\in{\mathbb{R}^{l_{f}\times{m}}}$, respectively.
In the Transformer architecture[43], the self-attention mechanism utilizes the learnable matrix $\mathbf{W}^{QKV}$ to generate the query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$) representations from inputs. Then, the attention is given by the scaled dot-product of $\mathbf{Q}$ and $\mathbf{K}$ with softmax normalization to achieve weighted information aggregation, which can effectively capture dependencies between different words in the sequence. Inspired by this, we input each $\mathbf{O}_{i}$ into stacked Transformer blocks to facilitate the interaction of the same object features in different frames in the $i$-th clip. Here, the number of frames in the clip corresponds to the sequence length in the original Transformer. By using self-attention mechanism, we can obtain an aggregated representation of each object $\mathbf{O}_{i,j,k}\in\mathbb{R}^{m}$ $(k\in\left[1,r\right])$ to learn the object-level temporal correlations. The aggregation can be expressed as:
@@
where $\alpha_{i,k,j,p}$ is the attention score. Eq. (7) denotes each object $\mathbf{O}_{i,j,k}$ is the fusion of the same object in different frames ($\mathbf{O}_{i,p,k},p\in\left[1,l_{f}\right]$) along the time dimension.
Subsequently, following the work [37], we construct a graph $\mathcal{G}_{i}$ for every $\mathbf{O}_{i,j}$, where each node and each edge in $\mathcal{G}_{i}$ denotes one object ($\mathbf{O}_{i,j,k}$) and the spatial correlations between two objects, respectively. Based on the constructed graph, we perform graph convolution operations on every $\mathbf{O}_{i,j}$, exploiting the spatial information between different objects in one frame to utilize spatial interaction. Finally, after average pooling along the object dimension, the processed object-level feature $\mathbf{O^{\prime}}_{i}\in\mathbb{R}^{l_{f}\times{m}}$ is concatenated with the frame-level feature $\mathbf{F}_{i}$. The concatenated feature is fed into a linear layer to map to the video semantics $\mathbf{Y}_{v,i}\in\mathbb{R}^{l_{f}\times{d}}$. All clip-level representations $\mathbf{Y}_{v,i}$ are concatenated to obtain the entire video semantics $\mathbf{Y}_{v}\in\mathbb{R}^{l_{v}\times{d}}$.
@@
III-B Cross-Attention Based DJSCC Transmission
@@
We use Transformer to construct the backbone of both the JSC encoder and decoder. In DJSCC, code rate guidance allows the refinement of the latent representations, thereby generating channel input symbols adaptive to bandwidth constraints, which motivates us to provide code rate guidance to the encoding and decoding processes of video semantics.
Different from directly using the human-defined code rate, we design the code rate as learnable parameters and interact with the video semantics in the form of cross attention. Specifically, we introduce a learnable parametric tensor, with the same shape as $\mathbf{Y}_{v}$, termed rate embedding $\mathbf{Y}_{\text{rate}}$, in the JSC encoder $f_{e}(\cdot;\boldsymbol{\theta},\boldsymbol{\epsilon})$ and the JSC decoder $f_{d}(\cdot;\boldsymbol{\phi})$. The encoding and decoding processes of $\mathbf{Y}_{v}$ are both guided by $\mathbf{Y}_{\text{rate}}$ shared between the transmitter and receiver. $\mathbf{Y}_{\text{rate}}$ allocates different attention to $\mathbf{Y}_{v}$ under different bandwidth constraints to aid encoding and decoding processes. Different from using random initialization, we set the initial values of $\mathbf{Y}_{\text{rate}}$ as a set of learnable parameters. By training under different bandwidth constraints, the model learns different initial values for $\mathbf{Y}_{\text{rate}}$. Each set of initial values can be viewed as prior information, representing the initial attention allocation results for semantics of different frames under a specific bandwidth constraint. During testing, for a specific bandwidth constraint, the corresponding initial values are selected to initialize $\mathbf{Y}_{\text{rate}}$ for inference. Furthermore, $\mathbf{Y}_{\text{rate}}$ can aid the JSC encoder in achieving variable-length coding of $\mathbf{Y}_{v}$, which will be explained in detail in Section III-C.
@@
$\tilde{\mathbf{Y}}_{\text{rate}}^{l}&=\operatorname{MHCA}(\operatorname{LN}(\mathbf{Y}_{\text{rate}}^{l-1}),\operatorname{LN}(\mathbf{Y}_{v}^{l-1}))+\mathbf{Y}_{\text{rate}}^{l-1},\\ \mathbf{Y}_{\text{rate}}^{l}&=\operatorname{FFN}(\operatorname{LN}(\tilde{\mathbf{Y}}_{\text{rate}}^{l}))+\operatorname{LN}(\tilde{\mathbf{Y}}_{\text{rate}}^{l}),\end{aligned}$ (10)
@@
$\tilde{\mathbf{Y}}_{v}^{l}&=\operatorname{MHCA}(\operatorname{LN}(\mathbf{Y}_{v}^{l-1}),\operatorname{LN}(\mathbf{Y}_{\text{rate}}^{l-1}))+\mathbf{Y}_{v}^{l-1},\\ \mathbf{Y}_{v}^{l}&=\operatorname{FFN}(\operatorname{LN}(\tilde{\mathbf{Y}}_{v}^{l}))+\operatorname{LN}(\tilde{\mathbf{Y}}_{v}^{l}),\end{aligned}$ (11)
respectively. In Eq. (10) and (11), $\mathbf{Y}_{\text{rate}}^{l-1}$ and $\mathbf{Y}_{v}^{l-1}$ denote the inputs of the $l$-th Transformer block of the two branches, and $\mathbf{Y}_{\text{rate}}^{l}$ and $\mathbf{Y}_{v}^{l}$ denote the outputs of the $l$-th Transformer blocks of the two branches. $\operatorname{MHCA}(\cdot)$ represents $\operatorname{CA}(\cdot)$ function with multi-heads. $\operatorname{LN}(\cdot)$ represents the layer normalization in the Transformer. $\operatorname{FFN}(\cdot)$ represents two linear layers with $\operatorname{GeLU}(\cdot)$ activation function.
@@
Figure 5: The JSC encoder with adaptive bandwidth allocation. CA Transformer block denotes the proposed dual-branch cross-attention Transformer block in Section III-B.
III-C Learning-Based Adaptive Bandwidth Allocation
@@
According to the cross-attention mechanism, $\mathbf{Y}_{\text{rate}}$ described in Section III-B can be seen as the score metric that dynamically scales elements of $\mathbf{Y}_{v}$, causing us to measure the importance of features based on $\mathbf{Y}_{\text{rate}}$. Our learning-based adaptive bandwidth allocation approach exploits $\mathbf{Y}_{\text{rate}}$ as guidance, which sparsifies the channels of each token output by the JSC encoder to generate channel input symbols of variable lengths. Specifically, given the output of the last dual-branch Transformer block $\mathbf{Y}_{\text{rate}}\in{\mathbb{R}^{l_{v}\times d}}$ and $\mathbf{Y}_{v}\in{\mathbb{R}^{l_{v}\times d}}$, we develop a series of rate predictors parameterized by $\boldsymbol{\epsilon}$ to predict the retained dimension for each token $\mathbf{Y}_{v,i}\in{\mathbb{R}^{d}}$ by $\mathbf{Y}_{\text{rate}}$. Then, a binary mask matrix $\mathbf{M}\in{\mathbb{R}^{l_{v}\times d}}$ is generated for channel masking, with the $i$-th row containing the first $k_{i}$ ones followed by $(d-k_{i})$ zeros ($k_{i}\in{\left[0,d\right]}$). $\mathbf{M}$ is used to retain the first $k_{i}$ channels of $\mathbf{Y}_{v,i}$ and mask the rest channels:
@@
where $\operatorname{AveragePool}(\cdot)$ denotes the average pooling operation. Then, $\mathbf{Z}^{l}_{\text{global}}$ is expanded from $\mathbb{R}^{1\times{d}}$ to $\mathbb{R}^{{l_{v}}\times{d}}$. After that, the rate predictor combines the local and global information along the channel dimension:
@@
At the receiver, we first generate $\mathbf{M}$ based on received $\mathbf{b}$. Then, we unflatten and zero-pad noisy real channel output symbols $\hat{\mathbf{s}}_{v}$ based on $\mathbf{M}$ to get $\hat{\mathbf{S}}_{v}$. The learnable vector $\mathbf{c}\in\mathbb{R}^{d}$ is developed to compensate for the information loss due to the channel masking operation. For each token $\hat{\mathbf{S}}_{v,i}\in\mathbb{R}^{d}$, if its $j$-th channel is masked, $\mathbf{c}_{j}$ is selected as the initial value for this channel:
@@
III-D Content-Adaptive and SNR-Adaptive
The bandwidth allocation method described in Section III-C can achieve bandwidth efficiency for VideoQA-SC. The decision $\mathbf{D}$ is only determined by a series versions of $\mathbf{Y}_{\text{rate}}$ ($\mathbf{Y}_{\text{rate}}^{1}$,$\mathbf{Y}_{\text{rate}}^{2}$,…,$\mathbf{Y}_{\text{rate}}^{L}$) and a series versions of $\mathbf{Y}_{v}$ ($\mathbf{Y}_{v}^{1}$,$\mathbf{Y}_{v}^{2}$,…,$\mathbf{Y}_{v}^{L-1}$), which indicates that the bandwidth allocation is adaptive to video semantics or video contents. Such rate predictors will make the same bandwidth allocation under different channel conditions, which is inconsistent with traditional channel coding ideas. Since content-adaptive bandwidth allocation is not robust to noise, it is difficult to support SC under diverse channel conditions.
@@ III-E Multimodal Fuser
Given $\hat{\mathbf{Y}}_{v}$ and $\mathbf{Y}_{q}$, the interaction of the two modal information can be achieved through attention-based weighted fusion. After linear projection, $\hat{\mathbf{Y}}_{v}$ is mapped to $\mathbf{E}_{v}\in\mathbb{R}^{l_{v}\times d}$ as query and $\mathbf{Y}_{q}$ is mapped to $\mathbf{E}_{q}\in\mathbb{R}^{b\times{s_{i}}\times d}$ as key. Then, we calculate the attention of $\hat{\mathbf{Y}}_{v}$ to each candidate QA pair feature $\mathbf{Y}_{q,i}$, and utilize attention-based fusion of $\hat{\mathbf{Y}}_{v}$ and $\mathbf{Y}_{q}$ to get the QA-aware video feature $\mathbf{Y}_{qv}\in\mathbb{R}^{l_{v}\times d}$, which can be formulated as:
@@
$a^{\star}=\mathop{\arg\max}\limits_{i}{\mathbf{a}}_{\text{pred},i},$ (27)
where ${\mathbf{a}}_{\text{pred},i}$ is the $i$-th element of $\mathbf{a}_{\text{pred}}$.
III-F Training Strategy